Find inverse laplace transform

(2s-56)/(s^2-4s-12)


1
Expert's answer
2021-11-18T15:31:20-0500

"\frac{2s-56}{s^2-4s-12}=\frac{2s-56}{(s+2)(s-6)}=\frac{15}{2(s+2)}-\frac{11}{2(s-6)}"

"L^{-1}(\frac{15}{2(s+2)}-\frac{11}{2(s-6)})=\frac{15}{2}e^{-2t}-\frac{11}{2}e^{6t}."